OpenStudy (anonymous):

help please. Evaluate. (2+i)/2-i

OpenStudy (anonymous):

multiply by (2+i)/(2+i)

OpenStudy (anonymous):

(2-i)(2+i)=4-i^2=5

OpenStudy (anonymous):

i don't get how it equals 5?

OpenStudy (anonymous):

the expression is not equal to 5; the denominator is. (2+i)(2+i)/((2-i)(2+i))=(2+i)^2/5=(3+4i)/5

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

hopefully that helps :)

OpenStudy (anonymous):

thank you!